1. A beam of protons, each with charge π‘ž β‰… 1.60 Γ— 10βˆ’19𝐢, is moving at 4.50 Γ— 105π‘š/𝑠 through a uniform magnetic


field with magnitude 0.60 T. The direction of motion of the protons is to the right, and the magnetic field direction is downward-right, at an angle of πœƒ = πœ‹ 6 π‘Ÿπ‘Žπ‘‘π‘–π‘Žπ‘›π‘  from the proton direction. Find the magnetic force acting on each


proton.




2. A straight copper wire carries a 6.0 A current from left to right in a region between the poles of large


electromagnet where a horizontal magnetic field directed at northeast with a magnitude of 2.10 T at 30 degrees


northeast. What is the magnetic force on a 2m section of a wire?

1
Expert's answer
2022-03-21T12:46:29-0400

(a)Magnetic force

"F=qvBsin\theta"

Put value


"F=1.6\times10^{-19}\times4.50\times10^5\times0.60\times sin30°"

"F=2.16\times10^{-14}N"

(B)

Magnetic force

"F=IlBsin\theta"


"F=6\times2\times2.10\times sin30°=12.6N"